John tosses a coin three times in succession. He wins $1 for head on the first toss, $2 for head on the second toss and $3 for head on the third toss. For each tail he losses $1.5. Find John's expected winnings. I saw the problem where the expected value was 81.25 cents but that is not one of my available answers which are 0.75, 1, 0.68 and -0.75.

Answer by stanbon(75887)

John tosses a coin three times in succession. He wins $1 for head on the first toss, $2 for head on the second toss and $3 for head on the third toss. For each tail he losses $1.5. Find John's expected winnings.
Head on 1st toss: (1/2)*1 = 0.5
Head on 2nd toss: (1/2)(-1.5)+(1/2)2 = -0.75+1 = 0.25
Head on 3rd toss: (1/2)(-1.5)+(1/2)(-1.5)+(1/2)3 = -0.75-0.75+1.5 = 0
-------------------------
Expected value = 0.75
---------------------------
Cheers,
Stan H.
===========================================
I saw the problem where the expected value was 81.25 cents but that is not one of my available answers which are 0.75, 1, 0.68 and -0.75.
---